OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively.

OpenStudy (anonymous):

@welshfella this is my last question

OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively. an = 9 • 4^n an = 9 • (-4)^(n + 1) an = 9 • 4^(n - 1) an = 9 • (-4)^(n - 1)

OpenStudy (anonymous):

From an example, I think you do something like \[2304=-36*r^3 \] or something like that

OpenStudy (welshfella):

second term is a1r and 5th term is ar^4 so r^4 / r = r^3 = 2304 / -36

OpenStudy (welshfella):

so r^3 = -64 so what is the value of r ?

OpenStudy (welshfella):

cube root of -64?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

So that's the first term?

OpenStudy (welshfella):

no that;s the common ratio r

OpenStudy (welshfella):

so thats excludes the first and third option where r = 4

OpenStudy (anonymous):

okay so we need to figure if it's n+ 1 or n-1

OpenStudy (anonymous):

when u finish welsh

OpenStudy (welshfella):

first term = a1 and second term = a1 * -4 = -36 so a1= -36 / -4 = ?

OpenStudy (anonymous):

9

OpenStudy (welshfella):

do u follow that?

OpenStudy (welshfella):

yes its 9 and the exponent is always n-1

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

is there anytime when it would be n+1?

OpenStudy (welshfella):

no

OpenStudy (anonymous):

welsh i respond the question u ask me on my post come take a look

OpenStudy (anonymous):

Okay thanks for all the help again! :)

OpenStudy (welshfella):

yw the general formula for the nth term an = a1. r^(n-1) - always n-1